OpenStudy (anonymous):

Find b and c if (2, -3) is the vertex of f(x) = 2x^2 + bx + c

jhonyy9 (jhonyy9):

f(2)=8+2b+c=0 ------ 2b+c= -8 f(-3)=2*9 -3b+c=0 --- -3b+c= -18 2b+c= -8 3b-c= 18 --------- 5b 0 = 10 b= 10/5 b=2 2*2+c= -8 4+c= -8 c= -8-4 = -12 c= -12 b=2 c=-12